Find the value of A and show work In(k) = In(A) + -E2(RT) k=0.00235 [a=-39.97 kJ R=8.314 J/mol*K T=273 K
Answer #1

- Criven that, R=0.00235 SR-81314J|molik Ea --39,97 KJ T=273K. Given equation es, en (K) = ln (A) + Ea(RT) en(0.00235) = N(A)